# CFG++: Manifold-constrained Classifier Free Guidance for Diffusion Models

### v1

@@ Abstract
Classifier-free guidance (CFG) is a fundamental tool in modern diffusion models for text-guided generation. Although effective, CFG has notable drawbacks. For instance, DDIM with CFG lacks invertibility, complicating image editing; furthermore, high guidance scales, essential for high-quality outputs, frequently result in issues like mode collapse. Contrary to the widespread belief that these are inherent limitations of diffusion models, this paper reveals that the problems actually stem from the off-manifold phenomenon associated with CFG, rather than the diffusion models themselves. More specifically, inspired by the recent advancements of diffusion model-based inverse problem solvers (DIS), we reformulate text-guidance as an inverse problem with a text-conditioned score matching loss, and develop CFG++, a novel approach that tackles the off-manifold challenges inherent in traditional CFG. CFG++ features a surprisingly simple fix to CFG, yet it offers significant improvements, including better sample quality for text-to-image generation, invertibility, smaller guidance scales, reduced mode collapse, etc. Furthermore, CFG++ enables seamless interpolation between unconditional and conditional sampling at lower guidance scales, consistently outperforming traditional CFG at all scales. Experimental results confirm that our method significantly enhances performance in text-to-image generation, DDIM inversion, editing, and solving inverse problems, suggesting a wide-ranging impact and potential applications in various fields that utilize text guidance. Project Page: https://cfgpp-diffusion.github.io/.
Figure 1: (Top) Comparison of T2I results by SDXL for the prompt "kayak in the water, optical color, aerial view, rainbow". The CFG-guided image has a broken oar, an artifact significantly reduced in the CFG++ version. (Middle) DDIM Inversion results under CFG show noticeable artifacts at various CFG scales, which are significantly reduced by CFG++. (Bottom) The evolution of denoised estimates differs between CFG and CFG++. CFG exhibits sudden shifts and intense color saturation early in reverse diffusion, while CFG++ transitions smoothly from low to high-resolution details.
@@ 1 Introduction
In modern text-to-image (T2I) diffusion models, the guidance scale $\omega$ is typically set within the range of $[5.0,30]$, referred to as the moderately high range of CFG guidance  (Chen et al., 2024b; Podell et al., 2023). The insufficiency in guidance also holds for classifier guidance (Dhariwal & Nichol, 2021; Song et al., 2021b) so that a scale of 10 was used. While using a high guidance scale yields higher-quality images with better alignment to the condition, it is also prone to mode collapse, reduces sample diversity, and yields an inevitable accumulation of errors during the sampling process. One example is DDIM inversion (Dhariwal & Nichol, 2021), a pivotal technique for controllable synthesis and editing (Mokady et al., 2023), where running the inversion process with $\omega>1.0$ leads to significant compromise in the reconstruction performance (Mokady et al., 2023; Wallace et al., 2023). Another extreme example would be score distillation sampling (SDS) (Poole et al., 2022), where the guidance scale in the order of a few hundred is chosen. Using such a high guidance scale leads to better asset quality to some extent, but induces blurry and saturated results. Several research efforts have been made to mitigate this downside by exploring methods where using a smaller guidance scale suffices (Wang et al., 2024; Liang et al., 2023). Although recent progress in SDS-type methods has reduced the necessary guidance scale to a range that is similar to those of ancestral samplers, using a moderately large $\omega$ is considered an inevitable choice.
@@
In this work, we aim to give an answer to this conundrum by revisiting the geometric view of diffusion models. In particular, inspired by the recent advances in diffusion-based inverse problem solvers (DIS) (Kadkhodaie & Simoncelli, 2021; Chung et al., 2023a; Song et al., 2023a; Kim et al., 2024; Chung et al., 2024), we reformulate the text-guidance as an inverse problem with a text-conditioned score matching loss and derive reverse diffusion sampling strategy by utilizing the decomposed diffusion sampling (DDS) (Chung et al., 2024). This results in a surprisingly simple fix of CFG to the sampling process without any computational overhead. The resulting process, which we call CFG++, works with a small guidance scale, typically $\lambda\in[0.0,1.0]$, that smoothly interpolates between unconditional and conditional sampling, with $\lambda=1.0$ having a similar effect as using CFG sampling with $\omega\sim 12.5$ at 50 neural function evaluation (NFE). Furthermore, DDIM inversion with CFG++ is invertible up to the discretization error, simplifying image editing. Comparing CFG++ against CFG shows that we achieve consistently better sample quality for text-to-image (T2I) generation, significantly better DDIM inversion capabilities that lead to enhanced reconstruction and editing, and enabling the incorporation of CFG guidance to diffusion inverse solvers (DIS) (Chung et al., 2023a). While the applications of CFG++ that we show in this work are limited, we believe that our work will have a broad impact that can be applied to all applications that leverage text guidance through the traditional CFG.
@@ 2 Background
Diffusion models. Diffusion models (Ho et al., 2020; Song et al., 2021b; Karras et al., 2022) are generative models designed to learn the reversal of a forward noising process. This process starts with an initial distribution $p_{0}({\boldsymbol{x}})$ where ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, and progresses towards the standard Gaussian distribution $p_{T}({\boldsymbol{x}})\approx\mathcal{N}(\bm{0},{\bm{I}})$, utilizing forward Gaussian perturbation kernels. Sampling from the data distribution can be performed by solving either the reverse stochastic differential equation (SDE) or the equivalent probability-flow ordinary differential equation (PF-ODE) (Song et al., 2021b). For example, under the choice $p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0})=\mathcal{N}({\boldsymbol{x}}_{0},t^{2}{\bm{I}})$, the generative PF-ODE reads
$d{\boldsymbol{x}}_{t}=-t\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})\,dt=\frac{{\boldsymbol{x}}_{t}-\mathbb{E}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]}{t}\,dt,\,\,{\boldsymbol{x}}_{T}\sim p_{T}({\boldsymbol{x}}_{T}),$ (1)
where Tweedie’s formula (Efron, 2011) ${\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]={\boldsymbol{x}}_{t}+t^{2}\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})$ is applied to achieve the second equality. Thus, the score function $\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})$ plays the key role in diffusion models. One can approximate an equivalent parameterization of the score function via denoising score matching (Hyvärinen & Dayan, 2005) or epsilon-matching (Ho et al., 2020), etc. When aiming for a text-conditional diffusion model that can condition on arbitrary ${\boldsymbol{c}}$, one can extend epsilon matching to a conditional one as:
$\min_{\theta}{\mathbb{E}}_{{\boldsymbol{x}}_{t}\sim p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0}),\{{\boldsymbol{x}}_{0},{\boldsymbol{c}}\}\sim p({\boldsymbol{x}}_{0},{\boldsymbol{c}}),{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\bm{I}})}\left[\|{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})-{\boldsymbol{\epsilon}}\|_{2}^{2}\right],$ (2)
where the condition is dropped with a certain probability (Ho & Salimans, 2021) so that null conditioning with ${\boldsymbol{c}}=\varnothing$ is possible. The neural network architecture of ${\boldsymbol{\epsilon}}_{\theta}$ is designed so that the condition ${\boldsymbol{c}}$ can modulate the output through cross attention (Rombach et al., 2022). For simplicity in notation throughout the paper, we define ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}:={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})$ and ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}:={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},\varnothing)$ by dropping $\theta$ and ${\boldsymbol{x}}_{t}$.
Extending the result of the Tweedie’s formula to the unconditional case under the variance preserving (VP) framework of DDPMs (Ho et al., 2020), we have
@@
where ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}:=\hat{\boldsymbol{x}}_{\varnothing}({\boldsymbol{x}}_{t})={\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t},\varnothing]$ is the denosed signal by Tweedie’s formula, and (5) corresponds to the renoising step. This is repeated for $t=T,T-1,\dots,1$.
@@
Diffusion model-based inverse problem solvers. Diffusion model-based inverse problem solvers (DIS) aim to perform posterior sampling from an unconditional diffusion model (kadkhodaie2021stochastic; chung2023diffusion; song2023pseudoinverseguided; kim2024dreamsampler). Specifically, for a given loss function $\ell({\boldsymbol{x}})$ which often stems from the likelihood for measurement consistency [^1] For example, for a given measurement model ${\boldsymbol{y}}=A{\boldsymbol{x}}+{\boldsymbol{w}}$ with ${\boldsymbol{w}}\sim\mathcal{N}({\boldsymbol{x}}_{0},\sigma^{2}{\bm{I}})$, the loss function is the data consistency (DC) loss given by $\ell({\boldsymbol{x}})=\|{\boldsymbol{y}}-A{\boldsymbol{x}}\|^{2}/2\sigma^{2}$. , the goal of DIS is to address the following optimization problem
$\min_{{\boldsymbol{x}}\in{\mathcal{M}}}\ell({\boldsymbol{x}})$ (8)
where ${\mathcal{M}}$ represents the clean data manifold sampled from unconditional distribution $p_{0}({\boldsymbol{x}})$. Consequently, it is essential to navigate in a way that minimizes cost while also identifying the correct clean manifold.
Recently, chung2023diffusion proposed a general technique called diffusion posterior sampling (DPS), where the updated estimate from the noisy sample ${\boldsymbol{x}}_{t}\in{\mathcal{M}}_{t}$ is constrained to stay on the same noisy manifold ${\mathcal{M}}_{t}$. This is achieved by computing the manifold constrained gradient (MCG) (chung2022improving) on a noisy sample ${\boldsymbol{x}}_{t}\in{\mathcal{M}}_{t}$ as $\nabla^{mcg}_{{\boldsymbol{x}}_{t}}\ell({\boldsymbol{x}}_{t}):=\nabla_{{\boldsymbol{x}}_{t}}\ell(\hat{\boldsymbol{x}}_{t})$, where $\hat{\boldsymbol{x}}_{t}$ is the denoised sample in (3) through Tweedie’s formula. The resulting algorithm can be stated as follows:
@@
where $\gamma_{t}>0$ denotes the step size. Under the linear manifold assumption (chung2022improving; chung2023diffusion), this allows precise transition to ${\mathcal{M}}_{t-1}$. Therefore, by performing (9) from $t=T$ to $t=0$, we can solve the optimization problem (8) with ${\boldsymbol{x}}_{0}\in{\mathcal{M}}$. Unfortunately, the computation of MCG requires computationally expensive backpropagation and is often unstable  (poole2022dreamfusion; du2023reduce).
In another work, chung2024decomposed shows that (9) can be equivalently represented as
${\boldsymbol{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}P_{\mathcal{M}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}},$ (10)
where $P_{\mathcal{M}}$ denotes the projection on the clean manifold ${\mathcal{M}}$. This implies that the data consistency update step becomes a standard gradient step, followed by the projection on the clean manifold. Importantly, under the commonly used linear manifold assumption in DPS, the one step update by ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})$ or multi-step update employing the conjugate gradient (CG) method are guaranteed to remain within a linear subspace, thus obviating the need for explicit computation of $P_{\mathcal{M}}$, i.e.
${\boldsymbol{x}}_{t-1}$ $\simeq\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}.$ (11)
This method, often referred to as the decomposed diffusion sampling (DDS), bypasses the computation of the score Jacobian, similar to poole2022dreamfusion, making it stable and suitable for large scale medical imaging inverse problems (chung2024decomposed). In the following, we leverage the insight from DDS to propose an algorithm to improve upon the CFG algorithm.

### v2

@@ Abstract
Classifier-free guidance (CFG) is a fundamental tool in modern diffusion models for text-guided generation. Although effective, CFG has notable drawbacks. For instance, DDIM with CFG lacks invertibility, complicating image editing; furthermore, high guidance scales, essential for high-quality outputs, frequently result in issues like mode collapse. Contrary to the widespread belief that these are inherent limitations of diffusion models, this paper reveals that the problems actually stem from the off-manifold phenomenon associated with CFG, rather than the diffusion models themselves. More specifically, inspired by the recent advancements of diffusion model-based inverse problem solvers (DIS), we reformulate text-guidance as an inverse problem with a text-conditioned score matching loss and develop CFG++, a novel approach that tackles the off-manifold challenges inherent in traditional CFG. CFG++ features a surprisingly simple fix to CFG, yet it offers significant improvements, including better sample quality for text-to-image generation, invertibility, smaller guidance scales, reduced mode collapse, etc. Furthermore, CFG++ enables seamless interpolation between unconditional and conditional sampling at lower guidance scales, consistently outperforming traditional CFG at all scales. Moreover, CFG++ can be easily integrated into the high-order diffusion solvers and naturally extends to distilled diffusion models. Experimental results confirm that our method significantly enhances performance in text-to-image generation, DDIM inversion, editing, and solving inverse problems, suggesting a wide-ranging impact and potential applications in various fields that utilize text guidance. Project Page: https://cfgpp-diffusion.github.io/.
Figure 1: (Top) Comparison of T2I results by SDXL-Lightning for the prompt "kayak in the water, optical color, aerial view, rainbow". The CFG-guided image has significant artifacts, which are reduced in the CFG++ version. (Middle) DDIM Inversion results under CFG show noticeable artifacts at various CFG scales, which are significantly reduced by CFG++. (Bottom) The evolution of denoised estimates differs between CFG and CFG++. CFG exhibits sudden shifts and intense color saturation early in reverse diffusion, while CFG++ transitions smoothly from low to high-resolution.
@@ 1 Introduction
In modern text-to-image (T2I) diffusion models, the guidance scale $\omega$ is typically set within the range of $[5.0,30]$, referred to as the moderately high range of CFG guidance  (Chen et al., 2024; Podell et al., 2023). The insufficiency in guidance also holds for classifier guidance (Dhariwal & Nichol, 2021; Song et al., 2021b) so that a scale of 10 was used. While using a high guidance scale yields higher-quality images with better alignment to the condition, it is also prone to mode collapse, reduces sample diversity, and yields an inevitable accumulation of errors during the sampling process. One example is DDIM inversion (Dhariwal & Nichol, 2021), a pivotal technique for controllable synthesis and editing (Mokady et al., 2023), where running the inversion process with $\omega>1.0$ leads to significant compromise in the reconstruction performance (Mokady et al., 2023; Wallace et al., 2023). Another extreme example would be score distillation sampling (SDS) (Poole et al., 2022), where the guidance scale in the order of a few hundred is chosen. Using such a high guidance scale leads to better asset quality to some extent, but induces blurry and saturated results. Several research efforts have been made to mitigate this downside by exploring methods where using a smaller guidance scale suffices (Wang et al., 2024; Liang et al., 2023). Although recent progress in SDS-type methods has reduced the necessary guidance scale to a range that is similar to those of ancestral samplers, using a moderately large $\omega$ is considered an inevitable choice.
@@
In this work, we aim to give an answer to this conundrum by revisiting the geometric view of diffusion models. In particular, inspired by the recent advances in diffusion-based inverse problem solvers (DIS) (Kadkhodaie & Simoncelli, 2021; Chung et al., 2023a; Song et al., 2023; Kim et al., 2024; Chung et al., 2024), we reformulate the text guidance as an inverse problem with a text-conditioned score-matching loss and derive a reverse diffusion sampling strategy by utilizing decomposed diffusion sampling (DDS) (Chung et al., 2024). This results in a surprisingly simple fix of CFG to the sampling process without any computational overhead. The resulting process, which we call CFG++, works with a small guidance scale, typically $\lambda\in[0.0,1.0]$, that smoothly interpolates between unconditional and conditional sampling, with $\lambda=1.0$ having a similar effect as using CFG sampling with $\omega\sim 12.5$ at 50 neural function evaluation (NFE). Furthermore, DDIM inversion with CFG++ is invertible up to the discretization error, simplifying image editing. Comparing CFG++ against CFG shows that we achieve consistently better sample quality for text-to-image (T2I) generation, significantly better DDIM inversion capabilities that lead to enhanced reconstruction and editing, and enabling the incorporation of CFG guidance to diffusion inverse solvers (DIS) (Chung et al., 2023a). While the applications of CFG++ that we show in this work are limited, we believe that our work will have a broad impact that can be applied to all applications that leverage text guidance through the traditional CFG.
@@ 2 Background
Diffusion models. Diffusion models (Ho et al., 2020; Song et al., 2021b; Karras et al., 2022) are generative models designed to learn the reversal of a forward noising process. This process starts with an initial distribution $p_{0}({\boldsymbol{x}})$ where ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, and progresses towards the standard Gaussian distribution $p_{T}({\boldsymbol{x}})\approx\mathcal{N}(\bm{0},{\bm{I}})$, utilizing forward Gaussian perturbation kernels. Sampling from the data distribution can be performed by solving either the reverse stochastic differential equation (SDE) or the equivalent probability-flow ordinary differential equation (PF-ODE) (Song et al., 2021b). For example, under the choice $p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0})=\mathcal{N}({\boldsymbol{x}}_{0},\sigma_{t}^{2}{\bm{I}})$, the generative PF-ODE reads
$d{\boldsymbol{x}}_{t}=-\sigma_{t}\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})\,dt=\frac{{\boldsymbol{x}}_{t}-\mathbb{E}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]}{\sigma_{t}}\,dt,\,\,\quad{\boldsymbol{x}}_{T}\sim p_{T}({\boldsymbol{x}}_{T}),$ (1)
where Tweedie’s formula (Efron, 2011) ${\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]={\boldsymbol{x}}_{t}+\sigma_{t}^{2}\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})$ is applied to achieve the second equality. When aiming for a text-conditional diffusion model that can condition on arbitrary ${\boldsymbol{c}}$, one can extend epsilon matching to a conditional one where the condition is dropped with a certain probability (Ho & Salimans, 2021) so that null conditioning with ${\boldsymbol{c}}=\varnothing$ is possible. The neural network architecture of ${\boldsymbol{\epsilon}}_{\theta}$ is designed so that the condition ${\boldsymbol{c}}$ can modulate the output through cross attention (Rombach et al., 2022). For simplicity in notation throughout the paper, we define ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}:={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})$ and ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}:={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},\varnothing)$ by dropping $\theta$ and ${\boldsymbol{x}}_{t}$.
Extending the result of Tweedie’s formula to the unconditional case under the variance preserving (VP) framework of DDPMs (Ho et al., 2020), we have
@@
where ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}:=\hat{\boldsymbol{x}}_{\varnothing}({\boldsymbol{x}}_{t})={\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t},\varnothing]$ is the denoised signal by Tweedie’s formula, and (4) corresponds to the renoising step. This is repeated for $t=T,T-1,\dots,1$.
@@
Diffusion model-based inverse problem solvers. Diffusion model-based inverse problem solvers (DIS) aims to perform posterior sampling from an unconditional diffusion model (kadkhodaie2021stochastic; chung2023diffusion; song2023pseudoinverseguided; kim2024dreamsampler). Specifically, for a given loss function $\ell({\boldsymbol{x}})$ which often stems from the likelihood, the goal of DIS is to address the optimization problem $\min_{{\boldsymbol{x}}\in{\mathcal{M}}}\ell({\boldsymbol{x}})$, where ${\mathcal{M}}$ represents the clean data manifold sampled from the unconditional distribution $p_{0}({\boldsymbol{x}})$. It is essential to navigate in a way that minimizes cost while also identifying the correct clean manifold.
chung2023diffusion proposed diffusion posterior sampling (DPS), where the updated estimate from the noisy sample ${\boldsymbol{x}}_{t}\in{\mathcal{M}}_{t}$ is constrained to stay on the same noisy manifold ${\mathcal{M}}_{t}$. This is achieved by computing the manifold constrained gradient (MCG) (chung2022improving) on a noisy sample ${\boldsymbol{x}}_{t}\in{\mathcal{M}}_{t}$ as $\nabla^{mcg}_{{\boldsymbol{x}}_{t}}\ell({\boldsymbol{x}}_{t}):=\nabla_{{\boldsymbol{x}}_{t}}\ell(\hat{\boldsymbol{x}}_{t})$, where $\hat{\boldsymbol{x}}_{t}$ is the denoised sample in (2) through Tweedie’s formula (efron2011tweedie). The resulting algorithm with DDIM (song2020score) can be stated as follows:
@@
where $\gamma_{t}>0$ denotes the step size. Under the linear manifold assumption (chung2022improving; chung2023diffusion), this allows precise transition to ${\mathcal{M}}_{t-1}$. To mitigate the computational complexity and the instability of neural network backprop, chung2024decomposed shows that (7) can be equivalently represented as
${\boldsymbol{x}}_{t-1}$ $\simeq\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ (8)
under further assumptions on ${\mathcal{M}}$. This method, often referred to as the decomposed diffusion sampling (DDS), bypasses the computation of the score Jacobian, similar to poole2022dreamfusion, making it stable and suitable for large-scale medical imaging inverse problems (chung2024decomposed). In the following, we leverage the insight from DDS to propose an algorithm to improve upon the CFG algorithm.